OpenStudy (anonymous):

given the linear equation y=-3/4 x-3, find the y-coordinates of the points (-8),(-4),(4).

OpenStudy (anonymous):

Just plug in those numbers for x and solve for y

OpenStudy (anonymous):

wolframalpha.com

OpenStudy (anonymous):

For example -8 y=-3/4* (-8)-3 y= 24/4 - 3 y= 6 -3 y= 3

OpenStudy (anonymous):

@enocrux Never use wolfram unless you need to check your work. Or you are doing a long integral.

OpenStudy (anonymous):

Now do it for the rest!

OpenStudy (anonymous):

it's a algebraic calculator, i have matlab and ti-89 for em, but for people who don't have calcs it's a help

OpenStudy (anonymous):

so Romero is that answer right

OpenStudy (anonymous):

For one of them. You have to do the rest. I'm not doing all of your homework

OpenStudy (anonymous):

you have to do -4 and 4

OpenStudy (anonymous):

i got the answer thanks

OpenStudy (anonymous):

Can you tell me what you got? I can check if they are right.

OpenStudy (anonymous):

y=-3/4x-3(-8),-4,4 (y=24 -3x/4, -4,4 -4=-4,4=4